# Conditional Feature Importance revisited: Double Robustness, Efficiency and Inference

### v4

@@ 3.2 Double robustness
We have observed that there is a need to estimate two regressors for both LOCO ($\widehat{m}$ and $\widehat{m}_{-j}$) and CPI ($\widehat{m}$ and $\widehat{\nu}_{-j}$). More generally, for CPI, we could use any conditional sampler $\widetilde{X}^{\prime(j)}$. In this section, we prove a double robustness property for detecting conditionally null covariates: to identify a null covariate, it is sufficient that one of the two estimates is consistent. This contrasts with LOCO, where errors in both estimates must compensate. This property explains the good empirical results obtained by CPI for variable selection (Chamma et al., 2024a ; Chamma et al., 2024b ; Paillard et al., (2025)).
@@
We first assume that $m$ depends only on the conditionally dependent features, and vice versa.
Assumption 3.4 (Functional and conditional equivalence).
$X^{j}\perp\!\!\!\perp y\mid X^{-j}$ if and only if $m(X)$ is not a function of $X^{j}$.
This assumption limits feature contributions only to higher-order effects—for example, a feature that is conditionally dependent only through its variance and has no predictive power. As shown in Proposition 3.1 of Reyero-Lobo et al., (2025), it holds under general ML settings, such as the additive noise model (3.7) below. For our results, this assumption is unnecessary, since we only require that conditional dependence implies functional dependence, which always holds.
We require that the ML model does not depend asymptotically on the unimportant features:
@@
This is a pointwise convergence on the null features, which can be easily verified for any GLM since the estimated null coefficients tend to 0. It is satisfied under the representability assumption for the Lasso (Zhao and Yu,, 2006). Under standard assumptions on Random Forests, the splits are performed with high probability only along the important features (Scornet et al.,, 2015, Proposition 1). For kernel methods, convergence in the RKHS together with the reproducing property implies pointwise convergence, and thus asymptotic relevance. For Gaussian processes, variable selection can be achieved using dimension-specific scalings (Bhattacharya et al.,, 2014). Finally, there are also results for 1-norm penalized SVM (Zhu et al.,, 2003) and neural networks (Dinh and Ho,, 2020). In Section D.2 we show numerically that this property holds across a large benchmark of ML models.
In any case, we note that both assumptions are weak and natural. Since our goal is to study conditional independence using a model, we are implicitly summarizing the XXyy relationship through the first-order moment ($\mathbb{E}\left[Y\mid X\right]$), and implicitly assuming that the model relies only on the important features. Thus, both assumptions are simply formalizations of what is expected from the model and data distribution when using ML models to study conditional independence.
@@
Assume that $X^{j}\perp\!\!\!\perp y\mid X^{-j}$. Given a conditional sampler such that $\widetilde{X}^{\prime(j)}\overset{\mathcal{L}}{\to}\widetilde{X}^{(j)}$, then for any bounded and continuous loss $\ell$ and continuous $\widehat{m}$, $\widehat{\psi}_{\mathrm{CPI}}\to 0$ a.s. On the other hand, given Assumption 3.5, for any conditional sampler $\widetilde{X}^{\prime(j)}$ and continuous loss $\ell$, $\widehat{\psi}_{\mathrm{CPI}}\to 0$ a.s.
The first implication of the theorem extends Theorem 2 from König et al., (2021) by making explicit the estimation of the conditional distribution. The second relies on the model’s implicit variable selection: as the loss optimizer, the model will generally not assign importance to irrelevant features, under the standard assumptions ensuring consistency of the ML model. For unbounded losses, the result holds if the loss is clipped with a sufficiently large constant.
@@
$\left.+2\left(m(\widetilde{X})-\widehat{m}(\widetilde{X})))(\widehat{m}(\widetilde{X})-\widehat{m}(\widetilde{X}^{\prime})\right)\middle|\mathcal{D}_{\mathrm{train}}\right].$
In this expression, we can observe the mentioned double robustness: it is sufficient to have either an accurate $\widehat{m}$ or $\widehat{\nu}_{-j}$. Specifically, the first error term will vanish either because $\widehat{m}$ detected the $j$-th coordinate as not important (and therefore the function does not change due to it), or because $\widehat{\nu}_{-j}$ is close to $\nu_{-j}$, making $\widetilde{X}^{\prime}$ close to $\widetilde{X}$. The second term is treated similarly.
@@
Therefore, there is a bias due to the estimation of both regressors. There are two main reasons why this error does not cancel out. The first one is an optimization error, which is more harmful in complex models. Indeed, it corresponds to a difference of errors from different models that have been independently optimized. Due to the variability of the optimization process, there may be many different solutions and multiple local minima, each with different errors, i.e. given $\widehat{m}_{1}$ and $\widehat{m}_{2}$ two optimizations, we do not forcely have $\mathbb{E}\left[(m(X)-\widehat{m}_{1}(X))^{2}\right]=\mathbb{E}\left[(m(X)-\widehat{m}_{2}(X))^{2}\right]$.
@@
In summary, LOCO compares errors of two independently optimized models, which do not cancel because (i) they are optimized separately and (ii) their error distributions differ. In contrast, Proposition 3.8 shows that CPI’s main error term compares the effect of estimating $\widetilde{X}$ within the same optimized model.
@@
In the remainder of this section, we focus on a Gaussian linear setting, which helps build intuition for CPI’s advantages over LOCO.
@@
We assume that $X$ is Gaussian. Thus, there is an explicit form for the conditional distribution: $X_{j}\mid X_{-j}=x_{-j}\sim\mathcal{N}(\mu^{j}_{\mathrm{cond}},\Sigma^{j}_{\mathrm{cond}})$ with $\mu^{j}_{\mathrm{cond}}:=\mu_{j}+\Sigma_{j,-j}\Sigma^{-1}_{-j,-j}(x_{-j}-\mu_{-j})$ and $\Sigma^{j}_{\mathrm{cond}}:=\Sigma_{j,j}-\Sigma_{j,-j}\Sigma^{-1}_{-j,-j}\Sigma_{-j,j}$. Then, $\nu_{-j}(X^{-j}):=\mathbb{E}\left[X^{j}|X^{-j}\right]$ is a linear model.
As seen in Lemma I.2, $y\mid X^{-j}$ is also a linear model. Thus, $\widehat{m}$, $\widehat{m}_{-j}$, and $\widehat{\nu}_{-j}$ are linear. However, the bias decays linearly for LOCO and quadratically for CPI.
@@
Assuming $X^{j}\perp\!\!\!\perp y\mid X^{-j}$, linear model (3.10), Gaussian covariates, and $\widehat{\nu}_{-j}$ and $\widehat{\beta}$ trained in different samples, then $\mathbb{E}\left[\widehat{\psi}_{\mathrm{CPI}}(j)\right]=O(1/n_{\mathrm{train}}^{2}).$
@@
Therefore, in this linear setting the double robustness translates as a faster convergence rate.
@@ 4 Sobol-CPI
where $\widetilde{X}^{(j)}\sim X\mid X^{-j}$. We propose to compute $1/n_{\mathrm{cal}}\sum_{i=1}^{n_{\mathrm{cal}}}\widehat{m}\left(\widetilde{x}^{{}^{\prime}(j)}_{i}\right)$, where $\{\widetilde{x}^{{}^{\prime}(j)}_{i}\}_{i=1,\dots,n_{\mathrm{cal}}}$ are sampled from the estimated conditional distribution. This can be easily achieved using CPI’s conditional sampling, proven valid in Section 3.1. We thus only need to train a regressor $\widehat{\nu}_{-j}$ and add $n_{\mathrm{cal}}$ residuals to obtain the estimate.
This idea is not limited to regression settings. Indeed, as discussed in Williamson et al., (2023), most Bayes predictors are functions of the conditional expectation. For example, under the classical 0-1 loss, the Bayes classifier is given by $\mathbb{I}_{\mathbb{E}[y|X]\geq 0.5}$. In this case, for the restricted model, we propose $\mathbb{I}_{1/n_{\mathrm{cal}}\sum_{i=1}^{n_{\mathrm{cal}}}\widehat{m}\left(\widetilde{x}^{(j)}_{i}\right)\geq 0.5},$ rather than refitting another model for $\widehat{m}_{-j}$. Then, we define the general Sobol-CPI as this TSI plug-in combination of the conditional sampling total’s expectation estimate of $m_{-j}$:
@@
Given a coordinate $j$, $n_{\mathrm{cal}}$, a regressor $\widehat{m}$ of $y$ given $X$ and a test set $(X_{i},y_{i})_{i=1,\ldots n_{\mathrm{test}}}$ , Sobol-CPI is defined as
$\widehat{\psi}_{\mathrm{SCPI}}^{j}$ $=\frac{1}{n_{\mathrm{test}}}\sum_{i=1}^{n_{\mathrm{test}}}\ell\left(\frac{1}{n_{\mathrm{cal}}}\sum_{k=1}^{n_{\mathrm{cal}}}\widehat{m}(\widetilde{x}^{{}^{\prime}(j)}_{i,k}),y_{i}\right)-\ell\left(\widehat{m}(x_{i}),y_{i}\right),$
where the $j$-th coordinate of $\widetilde{x}^{{}^{\prime}(j)}_{i,k}$ is conditionally sampled, and the rest fixed to $x_{i}^{-j}$ (see (1)).
We observe that this idea of using the law of total expectation is related to the marginalization of the conditional SAGE value functions (cSAGEvf, Covert et al., (2020)), while making explicit the fact that neither the conditional density nor the conditional expectation is known and needs to be estimated. Regarding the first issue, we study in particular the sampling step from the CPI and propose an asymptotic efficiency result. We further observe that, in order to extend this result to other cSAGEvf, it is necessary to specify the sampler explicitly. For the second issue, we propose a bias-correction procedure that yields an unbiased estimator without incurring additional cost.
@@ 4.1 Asymptotic efficiency
Under the assumptions of Williamson et al., (2023) (discussed in Appendix E) and an additional Lipschitz condition ensuring local robustness to small input changes, we achieve the same nonparametric efficiency.
@@
Among the assumptions, those on the loss function have been shown by Williamson et al., (2023) to hold for standard losses, such as the quadratic loss or classification accuracy. We also require the usual $O(n^{-1/4})$ convergence rate for $\widehat{m}$ and $\widehat{\nu}_{-j}$, which is standard in semiparametric inference and can be achieved by common ML models under mild conditions.
@@ 4.2 Fixing $n_{\mathrm{cal}}$ for the quadratic loss $\ell_{2}$
In practice, we need to decide on a finite value for $n_{\mathrm{cal}}$. Doing so introduces a trade-off between variable selection and variable importance. Indeed, as discussed in Section J.2, with a small $n_{\mathrm{cal}}$, we recover the double robustness of CPI, achieving better results for variable selection but losing nonparametric efficiency, leading to slightly worse results for variable importance. This is because the optimality assumption (Assumption A1 in Appendix E), which controls the first-order bias due to the need to estimate $m$, no longer holds. In any case, with the $\ell_{2}$-loss, fixing $n_{\mathrm{cal}}$ induces a bias:
@@
Given a coordinate $j$, a fixed $n_{\mathrm{cal}}$, a regressor $\widehat{m}$ of $y$ given $X$ and a test set $(X_{i},y_{i})_{i=1,\ldots n_{\mathrm{test}}}$ , Sobol-CPI($n_{\mathrm{cal}}$) is defined as
@@
where the $j$-th coordinate of $\widetilde{x}^{{}^{\prime}(j)}_{i,k}$ is conditionally sampled, and the rest fixed to $x_{i}^{-j}$ (see (1)).
@@
In particular, by taking $n_{\mathrm{cal}}=1$, we recover the standard CPI, but divided by 2. With this simple correction, the method works effectively for variable selection, as we directly recover double robustness. Moreover, it will converge to TSI, even though it is not efficient (see Section J.2). This result also corrects Theorem 2 from Hooker et al., (2021), which claimed that Dropped and Conditional Variable Importance coincide, highlighting the gap in the literature between the two approaches.
4.3 Confidence intervals
@@
As stated in Theorem 4.2, there is no need for a one-step estimate to correct any first-order bias. The same holds for LOCO, which does not require such corrections when estimated as a plug-in estimate in the difference of generalized ANOVA (Williamson et al., (2021)) or as a difference of predictiveness measures more generally (Williamson et al., (2023)). Whenever the importance is not null, it is possible to construct confidence intervals using the influence functions. Indeed, the variance is given by $0<\tau_{j}^{2}:=\mathbb{E}[\varphi_{j}^{2}(z)]<\infty$, where $\varphi_{j}$ is the influence function of $\psi_{\mathrm{TSI}}(j,P_{0})$. Therefore, it is possible to estimate it using a plug-in estimate (Williamson et al., (2023)).
@@
Moreover, as shown in Appendix H, when taking the MSE as the predictiveness measure, the variance estimated with this plug-in method matches the natural empirical variance exactly. Nevertheless, under the null hypothesis, the influence function vanishes, preventing a Gaussian asymptotic distribution. This poses a challenge, as it hinders variable selection with direct statistical guarantees—an essential component of reliable scientific discovery. This issue is the same as that faced by LOCO and Shapley due to the quadratic functionals (Verdinelli and Wasserman, (2024)).
@@
Several approaches have been proposed to address this problem. Williamson et al., (2023) noted that even if the influence function of the variable importance measure vanishes, excluding extreme cases, the influence functions of the predictiveness measures with and without the covariate do not vanish. They attempted to leverage this observation by computing each predictiveness measure on different data splits. However, as observed in numerical experiments in Section 5, this approach does not work well because it does not only increase variability but also significantly increases bias, resulting in a loss of power and poor estimates.
@@
Other alternatives include inflating the confidence interval by an additive term of the order $O_{P}(\sqrt{n})$ (Dai et al., (2024); Verdinelli and Wasserman, (2024)). Validity can then be obtained using Chebyshev’s inequality, resulting in a confidence interval of the form $(-\infty,\widehat{\psi}_{n}+z_{\alpha}\mathrm{se}_{n}+c/\sqrt{n}]$, where $\mathrm{se}_{n}$ is the empirical standard error and $c$ is any constant. In practice, $c$ is taken as the standard error of the output, similar to Verdinelli and Wasserman, (2024). We use the same correction idea.
@@
Using Theorem 4.2, we establish the consistency of this conditional null hypothesis test with the additive term (see Appendix G). Furthermore, based on Lemma 3.11, we propose applying Markov’s inequality to construct a more powerful test—achieving a rate of $c/n^{\gamma}$ for $\gamma<2$ instead of the standard $c/\sqrt{n}$—as demonstrated in Appendix G. The limitations regarding the extent of interval expansion are discussed in Appendix J.4, both for linear and nonlinear settings.

### v5

@@ 3.2 Double robustness
We have observed that there is a need to estimate two models for both LOCO ($\widehat{m}$ and $\widehat{m}_{-j}$) and CPI ($\widehat{m}$ and $P^{\prime}_{j}$). In this section, we prove a double robustness property for detecting conditionally null covariates: to identify a null covariate, it is sufficient that one of the two estimates is consistent. This contrasts with LOCO, where errors in both estimates must compensate. This property explains the good empirical results obtained by CPI for variable selection (Chamma et al., 2024a, b; Paillard et al., 2025).
@@
Note that under the null hypothesis, the theoretical model satisfies $m(X)=\mathbb{E}[y\mid X]=\mathbb{E}[y\mid X^{-j}]$ since $y\perp\!\!\!\perp X^{j}\mid X^{-j}$; hence, it does not depend on the null feature. We require that the ML model asymptotically does neither depend on unimportant features:
@@
This is a pointwise convergence on the null features, which can be easily verified for any GLM since the estimated null coefficients tend to 0. It is satisfied under the representability assumption for the Lasso (Zhao and Yu, 2006). Under standard assumptions on Random Forests, the splits are performed with high probability only along the important features (Prop. 1 Scornet et al. (2015)). For kernel methods, convergence in the RKHS together with the reproducing property implies pointwise convergence, and thus asymptotic relevance. For Gaussian processes, variable selection can be achieved using dimension-specific scalings (Bhattacharya et al., 2014). Finally, there are also results for 1-norm penalized SVM (Zhu et al., 2003) and neural networks (Dinh and Ho, 2020). In Sections D.2 and J.5.3, we show numerically that this property holds across a broad benchmark of ML models on both simulated and real datasets.
In any case, since our goal is to study conditional independence using a model, we are implicitly summarizing the $X$-$y$ relationship through the first-order moment ($\mathbb{E}\left[Y\mid X\right]$), and implicitly assuming that the model mainly relies only on the important features. Thus, the assumption is natural and is a formalization of what is expected from $\widehat{m}$ when using ML models to study conditional independence.
@@
Assume that $X^{j}\perp\!\!\!\perp y\mid X^{-j}$. Given a conditional sampler $P^{\prime}_{j}$ such that $\widetilde{X}^{\prime(j)}\overset{\mathcal{L}}{\to}\widetilde{X}^{(j)}$, then for any bounded and continuous loss $\ell$ and continuous $\widehat{m}$, $\widehat{\psi}_{\mathrm{CPI}}\to 0$ a.s. On the other hand, given Assumption 3.4, for any conditional sampler $P^{\prime}_{j}$ and continuous loss $\ell$, $\widehat{\psi}_{\mathrm{CPI}}\to 0$ a.s.
The first implication of the theorem extends Theorem 2 from König et al. (2021) by making explicit the estimation of the conditional distribution. The second relies on the model’s implicit variable selection: as the loss optimizer, the model will generally not assign importance to irrelevant features, under the standard assumptions ensuring consistency of the ML model. For unbounded losses, the result holds if the loss is clipped with a sufficiently large constant.
@@
$\left.+2\left((m(\widetilde{X})-\widehat{m}(\widetilde{X})))(\widehat{m}(\widetilde{X})-\widehat{m}(\widetilde{X}^{\prime})\right)\middle|\mathcal{D}_{\mathrm{train}}\right].$
In this expression, we can observe the mentioned double robustness: it is sufficient to have either an accurate $\widehat{m}$ or $P^{\prime}_{j}$. Specifically, the first error term will vanish either because $\widehat{m}$ detected the $j$-th coordinate as not important (and therefore the function does not change due to it), or because $P^{\prime}_{j}$ is close to $P^{\star}_{j}$, making $\widetilde{X}^{\prime}$ close to $\widetilde{X}$. The second term is treated similarly.
@@
Therefore, there is a bias due to the estimation of both regressors. There are two main reasons why this error does not cancel out. The first one is an optimization error, which is more harmful in complex models (Bottou and Bousquet (2007)). Indeed, it corresponds to a difference of errors from different models that have been independently optimized. Due to the variability of the optimization process, there may be many different solutions and multiple local minima, each with different errors, i.e. given $\widehat{m}_{1}$ and $\widehat{m}_{2}$ two optimizations, we do not forcely have $\mathbb{E}\left[(m(X)-\widehat{m}_{1}(X))^{2}\right]=\mathbb{E}\left[(m(X)-\widehat{m}_{2}(X))^{2}\right]$.
@@
In summary, LOCO compares errors of two independently optimized models over different sets of features, which do not cancel because (i) they are optimized separately and (ii) their error distributions differ. In contrast, Proposition 3.7 shows that CPI’s main error term compares the effect of estimating $\widetilde{X}^{(j)}$ within the same optimized model.
@@
In the remainder of this section, we focus on a Gaussian linear setting, which helps build intuition for CPI’s advantages over LOCO. Indeed, rather than requiring only one model to work to detect a null as before, we show that when both models are accurate, the null is detected faster. We use the regression-based sampler (1) to obtain explicit convergence rates.
@@
We assume that $X$ is Gaussian. Thus, there is an explicit form for the conditional distribution: $X_{j}\mid X_{-j}=x_{-j}\sim\mathcal{N}(\mu^{j}_{\mathrm{cond}},\Sigma^{j}_{\mathrm{cond}})$ with $\mu^{j}_{\mathrm{cond}}:=\mu_{j}+\Sigma_{j,-j}\Sigma^{-1}_{-j,-j}(x_{-j}-\mu_{-j})$ and $\Sigma^{j}_{\mathrm{cond}}:=\Sigma_{j,j}-\Sigma_{j,-j}\Sigma^{-1}_{-j,-j}\Sigma_{-j,j}$. In particular, $\nu_{-j}(X^{-j}):=\mathbb{E}\left[X^{j}|X^{-j}\right]$ is a linear model.
As shown in Lemma I.2, $y\mid X^{-j}$ is also linear. Consequently, $\widehat{m}$, $\widehat{m}_{-j}$, and $\widehat{\nu}_{-j}$ are linear, so LOCO and CPI have the same computational complexity. However, LOCO’s bias decays linearly, whereas CPI’s decays quadratically.
@@
Assuming $X^{j}\perp\!\!\!\perp y\mid X^{-j}$, linear model (3.9), Gaussian input covariates, and $\widehat{\nu}_{-j}$ and $\widehat{\beta}$ trained in different samples, then $\mathbb{E}\left[\widehat{\psi}_{\mathrm{CPI}}(j)\right]=O(1/n_{\mathrm{train}}^{2}).$
@@
Therefore, in this linear setting the double robustness translates as a faster convergence rate for CPI.
@@ 4 Sobol-CPI
where $\widetilde{X}^{(j)}\sim\mathcal{L}(X\mid X^{-j})$. We propose to compute $1/n_{\mathrm{cal}}\sum_{i=1}^{n_{\mathrm{cal}}}\widehat{m}\left(\widetilde{x}^{{}^{\prime}(j)}_{i}\right)$, where $\{\widetilde{x}^{{}^{\prime}(j)}_{i}\}_{i=1,\dots,n_{\mathrm{cal}}}$ are sampled from the estimated conditional distribution $P^{\prime}_{j}$. This can be easily achieved using CPI’s conditional sampling, proven valid in Section 3.1. We thus only need to train a regressor $\widehat{\nu}_{-j}$ and add $n_{\mathrm{cal}}$ residuals to obtain the estimate.
This idea is not limited to regression settings. Indeed, as discussed in Williamson et al. (2023), most Bayes predictors are functions of the conditional expectation. For example, under the classical 0-1 loss, the Bayes classifier is given by $\mathbb{I}_{\mathbb{E}[y|X]\geq 0.5}$. In this case, for the restricted model, we propose $\mathbb{I}_{1/n_{\mathrm{cal}}\sum_{i=1}^{n_{\mathrm{cal}}}\widehat{m}\left(\widetilde{x}^{(j)}_{i}\right)\geq 0.5},$ rather than refitting another model for $\widehat{m}_{-j}$. Then, we define the general Sobol-CPI as this TSI plug-in combination of the conditional sampling total’s expectation estimate of $m_{-j}$:
@@
Given a coordinate $j$, a calibration set size $n_{\mathrm{cal}}$, a regressor $\widehat{m}$ of $y$ given $X$ and a test set $\{(X_{i},y_{i})\}_{i=1,\ldots n_{\mathrm{test}}}$ , Sobol-CPI is defined as
$\hskip-5.69054pt\widehat{\psi}_{\mathrm{SCPI}}^{j}$ $=\frac{1}{n_{\mathrm{test}}}\sum_{i=1}^{n_{\mathrm{test}}}\ell\left(\frac{1}{n_{\mathrm{cal}}}\sum_{k=1}^{n_{\mathrm{cal}}}\widehat{m}(\widetilde{x}^{{}^{\prime}(j)}_{i,k}),y_{i}\right)-\ell\left(\widehat{m}(x_{i}),y_{i}\right)$
where the $j$-th coordinate of $\widetilde{x}^{\prime(j)}_{i,k}$ is sampled from an estimated conditional distribution $P^{\prime}_{j}$, with the remaining coordinates fixed to $x_{i}^{-j}$.
We observe that using the law of total expectation is related to the marginalization of the conditional SAGE value functions (cSAGEvf, Covert et al. (2020)), while making explicit the fact that both the conditional density and the conditional expectation are unknown and must estimated. Regarding the first issue, we study in particular the sampling step from the CPI and propose an asymptotic efficiency result. We further observe that, in order to extend this result to cSAGEvf, it is necessary to specify the sampler explicitly. For the second issue, we propose a bias-correction procedure that yields an unbiased estimator without incurring additional cost.
@@ 4.1 Asymptotic efficiency
Under the assumptions of Williamson et al. (2023) (discussed in Appendix E) and an additional Lipschitz condition ensuring local robustness to small input changes, we achieve the same Sobol-CPI’s nonparametric efficiency.
@@
Among the assumptions, those on the loss function have been shown by Williamson et al. (2023) to hold for standard losses, such as the quadratic loss or classification accuracy. We also require the usual $O(n^{-1/4})$ convergence rate for $\widehat{m}$ and $\widehat{\nu}_{-j}$, which is standard in semiparametric inference. Note that this result holds for any conditional sampler $P^{\prime}_{j}$ such that $\|\widetilde{X}^{\prime(j)}-\widetilde{X}^{(j)}\|=O_{P}(n^{-1/4})$.
@@ 4.2 Fixing $n_{\mathrm{cal}}$ for the quadratic loss $\ell_{2}$
In practice, we need to decide on a finite value for $n_{\mathrm{cal}}$. Doing so introduces a trade-off between variable selection and variable importance. Indeed, as discussed in Section J.2, with a small $n_{\mathrm{cal}}$, we recover the double robustness of CPI, achieving better results for variable selection but losing nonparametric efficiency, leading to slightly worse results for variable importance. This is because the optimality assumption (Assumption A1 in Appendix E), which controls the first-order bias due to the need to estimate $m$, no longer holds. In any case, fixing $n_{\mathrm{cal}}$, as is commonly done in practice for SAGE (Covert et al., 2021), induces a bias:
@@
Given a coordinate $j$, a fixed $n_{\mathrm{cal}}$, a regressor $\widehat{m}$ of $y$ given $X$ and a test set $\{(X_{i},y_{i})\}_{i=1,\ldots n_{\mathrm{test}}}$ , Sobol-CPI($n_{\mathrm{cal}}$) is defined as
@@
where the $j$-th coordinate of $\widetilde{x}^{\prime(j)}_{i,k}$ is sampled from an estimated conditional distribution $P^{\prime}_{j}$, with the remaining coordinates fixed to $x_{i}^{-j}$.
@@
In particular, by taking $n_{\mathrm{cal}}=1$, we recover the standard CPI, but divided by 2. With this simple correction, the method works effectively for variable selection, as we directly recover double robustness. Moreover, it converges to TSI, even though it is not efficient (see Section J.2). This result also corrects Theorem 2 from Hooker et al. (2021), which claimed that Dropped and Conditional Variable Importance coincide, highlighting the gap in the literature between the two approaches.
5 Confidence intervals and Inference
@@
As stated in Theorem 4.2, there is no need for a one-step estimate to correct any first-order bias. The same holds for LOCO, which does not require such corrections when estimated as a plug-in estimate in the difference of generalized ANOVA (Williamson et al., 2021) or as a difference of predictiveness measures more generally (Williamson et al., 2023). Whenever the importance is not null, it is possible to construct confidence intervals using the influence functions. Indeed, the variance is given by $0<\tau_{j}^{2}:=\mathbb{E}[\varphi_{j}^{2}(z)]<\infty$, where $\varphi_{j}$ is the influence function of $\psi_{\mathrm{TSI}}(j,P_{0})$. Therefore, it is possible to estimate it using a plug-in estimate. Moreover, as shown in Appendix H, when taking the MSE as the predictiveness measure, the variance estimated with this plug-in method matches the natural empirical variance.
@@
Nevertheless, under the null hypothesis, the influence function vanishes, preventing a Gaussian asymptotic distribution. This poses a challenge, as it hinders variable selection with direct statistical guarantees—an essential component of reliable scientific discovery. This issue is the same as that faced by LOCO and Shapley due to the quadratic functionals (Verdinelli and Wasserman, 2024).
@@
To address this problem, Williamson et al. (2023) noted that even if the influence function of LOCO vanishes, excluding extreme cases, the influence functions of the predictiveness measures with and without the covariate do not vanish. They attempted to leverage this observation by computing each predictiveness measure on different data splits. However, as observed in numerical experiments in Section 6, this is ineffective because it increases both variability and bias, resulting in a loss of power and poor estimates.
@@
Other alternatives include inflating the confidence interval by an additive term of the order $O_{P}(\sqrt{n})$ (Dai et al. (2024); Verdinelli and Wasserman (2024)). Validity can then be obtained using Chebyshev’s inequality, resulting in a confidence interval of the form $(-\infty,\widehat{\psi}_{n}+z_{\alpha}\mathrm{se}_{n}+c/\sqrt{n}]$, where $\mathrm{se}_{n}$ is the empirical standard error and $c$ is any constant. In practice, Verdinelli and Wasserman (2024) takes $c$ as the standard error of the output.
We first propose a similar variance correction improving the rates using our previous results, and then introduce a finite-sample Wilcoxon test that also accounts for the estimation of the conditional distribution.
Variance correction:
@@
Using Theorem 4.2, we establish the consistency (power tending to $1$) of this conditional hypothesis test with the additive correction (see Appendix G.1). Moreover, using Lemma 3.10, we propose applying Markov’s inequality to construct a more powerful test, achieving a rate of $c/n^{\gamma}$ for $\gamma<2$ instead of the standard $c/\sqrt{n}$ (Appendix G.1). The limitations regarding the extent of interval expansion are discussed in Appendix J.4, both for linear and nonlinear settings.
Nonparametric tests:
Variance-correction procedures rely on asymptotic convergence rates, which are typically not tight and do not provide finite-sample guarantees. To address this, in addition to the corrections motivated by our convergence results from previous sections, as detailed in Algorithm 1, we propose using a paired nonparametric test, such as the Wilcoxon or Sign test, between
$\left\{l\bigl(\widehat{m}(X_{i}),y_{i}\bigr)\right\}_{i}^{n_{\mathrm{test}}}\text{ and }\left\{l\bigl(\widehat{m}(\widetilde{X}^{\prime(j)}_{i}),y_{i}\bigr)\right\}_{i}^{n_{\mathrm{test}}}.$
This yields a p-value $p^{\prime}$. Any conditional sampler may be used; if the theoretical conditional was available, this would provide exact finite-sample type-I error. Since the conditional must be estimated in practice, we provide a bound on the error based on the Total Variation distance $d_{\mathrm{TV}}$ between the estimated and true conditional distributions. For more details on this distance, see e.g. Berrett et al. (2019).
Theorem 5.1.
Let $P^{\star}_{j}$ denote the conditional distribution $\mathcal{L}(X^{j}\mid X^{-j})$, estimated by any $P^{\prime}_{j}$. Then, for any level $\alpha\in[0,1]$, the p-value $p^{\prime}$ produced by Algorithm 1 satisfies
$\mathbb{P}_{\mathcal{H}_{0}}($ $p^{\prime}\leq\alpha\mid\{X^{-j}_{i},y_{i}\}_{i}^{n_{\mathrm{test}}},\mathcal{D}_{\mathrm{train}})$
$\leq\alpha+d_{\mathrm{TV}}(P^{\prime}_{j}(\cdot\mid X^{-j}),P^{\star}_{j}(\cdot\mid X^{-j})).$